Proceedings of AOSE: Camera-Ready Copy Preparation

Person One^{∗} Person Two^{†} Person Three^{‡}

∗ Mitsubishi Electric Digital Library Group, 17th Floor, Centre Point

103 New Oxford St, London WC1A 1EB, U.K.

         pa@dlib.com

† Department of Computing, Manchester Metropolitan University

Manchester M1 5GD, U.K.

         pb@doc.mmu.ac.uk

‡ Department of Electronic Engineering, Queen Mary and Westfield College

Mile End Road, London E1 4NS, U.K.

         pc@qmw.ac.uk

Abstract. The AOSE proceedings are formally published by Springer-Verlag in their Lecture Notes in AI series. The aim of this document is to give detailed formatting instructions for preparation of camera ready copy. Please take the time to read these instructions, even if you have prepared LNAI/LNCS articles before. This document is formatted in the correct way; it may help to use it as an model of what your article should look like.

1 Introduction

The proceedings of AOSE are published by Springer-Verlag in their Lecture Notes in AI (LNAI) series. LNAI volumes are published from camera ready copy prepared by individual authors. In this document, we set out detailed formatting requirements for AOSE authors.

Please take a few minutes to read it, even if you believe you are familiar with the preparation of articles for LNAI volumes.

Please note: We reserve the right not to publish any article that, in our judgement, has:

– not been revised according to the requirements set out in reviews; or

– not been prepared in accordance with the camera ready copy requirements set out here; or

– exceeded the page limit (see below); or

– not met the deadline for submission of camera ready copy.

We will be fussy about details such as formatting of references, scruffy layout, and so on! If in doubt, this document is an example of a correctly formatted article.

2 Main Points

Here are the main points that everyone will need to note:

Font: Your article should be set in PostScript times roman font. For L^{A}TEX users,

this means using the times package. Title: If the length of the title exceeds one line, then break the line, ensuring that the

resulting lines are more or less of the same length. Author names and contact details: Please give the name and full postal address of

every author, along with an email address if possible. Please do not (i) use initials instead of first names; or (ii) give phone numbers, fax numbers, or WWW addresses. If you have authors from different institutions, then look at this document to see how to format names and addresses; the file skeleton.tex provided in the author pack gives the desired format for single-authored papers.

Sections: Please do not put section names in UPPER CASE, AS IT LOOKS UGLY! The default rule is to capitalise the first letter of every word longer than five letters, and the first letter of every other ‘significant’ word (similarly for the title of the article). The format for section number is shown in this document. Please don’t have too many levels of section numbering — a section numbered 4.2.1 is guaranteed to send your reader to sleep.

References: L^{A}TEX users should use the plain bibliography style. This means that

references are referred to in the text using [18] notation. Royalties: There are no royalties (for anyone) from LNAI volumes!

2.1 L^{A}TEX Users

L^{A}TEX users will have an easy time of it: there is a class (llncs.cls — included in this author pack) which can be used to achieve most of the desired results. However, this style file does not do everything. You should ensure that:

– You typeset your article using the PostScript times (roman) font, not the Computer Modern default. (This is achieved by using the times style file, as indicated above.)

– You use the plain bibliography style (produces numbers in square brackets).

– Your lines or diagrams do not fall outside the 12.2 × 19.3cm printing area. This is very easy to miss when processing with L^{A}TEX, and Springer will not publish a volume that contains such lines. You will have to pay attention to those ‘Overfull hbox’ messages!

2.2 Non-L^{A}TEX Users

Microsoft Word users can use the style file sv-lncs.dot, included in this author pack. Macintosh users can use the sv-lncs file. If you use either of these, we ask that you pay particular attention to ensure that your final camera ready copy looks like the L^{A}TEX produced versions. Print out skeleton.ps and instructions.ps for details.